A cashier saves 1/5 of her gross monthly salary. This is $40 more than is saved by her friend that is a secretary whose salary is $200 greater. If the secretary save 1/7 of her salary. How much does each earn per month?
Answer by ankor@dixie-net.com(12704)

A cashier saves 1/5 of her gross monthly salary. This is $40 more than is saved by her friend that is a secretary whose salary is $200 greater. If the secretary save 1/7 of her salary. How much does each earn per month?
:
Let x = cashier's salary
Secretary earns $200 more so we can say her salary = (x+200)
:
Write an expression for each given statement:
:
"A cashier saves 1/5 of her gross monthly salary."
(1/5)x
:
"the secretary save 1/7 of her salary."
(1/7)(x+200)
:
"This is $40 more than is saved by her friend that is a secretary"
:
Cashier's saving = Secretary's savings + $40
(1/5)x = (1/7)(x+200) + 40
:
Multiply thru by 35 to get rid of the denominators, resulting in:
7x = 5(x+200) + 35(40)
7x = 5x + 1000 + 1400
7x - 5x = 2400
2x = 2400
x = 2400/2
x = $1200 is the cashier's salary, obviously $1400 is the secretary's salary
:
:
We can check this easily
(1/5)1200 = 240
(1/7)1400 = 200
---------------
difference: 40